The base of a triangle is 15 inches and its area is 60 square inches. Find the area of a similar triangle of altitude 3.
Answer by checkley77(12844)

AREA=BH/2
60=15H/2
15H=120
H=120/15
H=8 ANS FOR THE HEIGHT.
8/15=3/X
8X=45
X=45/8
X=5.625 ANS. FOR THE BASE OF THE SIMILIAR TRIANGLE.
PROOF:
8/15=3/5.625
15*3=8*5.625
45=45